OpenStudy (anonymous):

suppose you choose a marble from a bag containing 4 red marbles, 2 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find the P(red and blue).

OpenStudy (anonymous):

would it be 7/9

OpenStudy (anonymous):

There are two rules: The AND rule The OR rule

OpenStudy (anonymous):

In the AND rule you multiply but in the OR rule you add

OpenStudy (anonymous):

But since its saying P(red and blue) that means you will multiply not add

OpenStudy (anonymous):

ok so you add 4 times 2 times 3?

OpenStudy (anonymous):

These are independent events because you are doing this "with replacement". So, you would take the first probability multiplied by the second probability multiplied by 2 because you could have red and then blue or blue and then red. Start with P(red) for just one pick, can you determine that first?

OpenStudy (anonymous):

No, your answer will be P(red) x P(blue) x 2. Let's work on P(red) first. Nine marbles in all of which 4 are red. So, P(red) = (# red)/(total # of marbles)

OpenStudy (anonymous):

P(blue) is similar to P(red) but with different numbers. 3 blue marbles, but still 9 total marbles in all. How are you coming along with P(red)?

OpenStudy (anonymous):

so P(red)=(4)/(24)

OpenStudy (anonymous):

so P(red)= (4)/(9)

OpenStudy (anonymous):

Yes! Now do P(blue).

OpenStudy (anonymous):

P(Blue)=(3)/(9)

OpenStudy (anonymous):

Good! Now, as stated above P(red and blue) = P(red) x P(blue) x 2 because the problem reads like it allows for either red or blue to be picked first. That's why there is a factor of "2" added on.

OpenStudy (anonymous):

ok Im with you im just a little confused

OpenStudy (anonymous):

(4)/(9)*(3)/(9)* 2

OpenStudy (anonymous):

np. Once we have this overall view, we need to drill down into understanding. The whole key to this problem begins with "replacing", so we always have 9 total marbles, we also have independence (which tells us to multiply the probabilities), and we have a factor of "2" because the order doesn't matter. Ask questions now. Just saw your answer and it's right! Good job!

OpenStudy (anonymous):

ok so what would the answer be then because i have to change it to fraction form

OpenStudy (anonymous):

it would be 36*27*2?

OpenStudy (anonymous):

\[\frac{ 4 }{ 9 }\times \frac{ 3 }{ 9 }\times2 = \frac{ 8 }{ 27 }\]

OpenStudy (anonymous):

but 8/27 is not on my answer thing

OpenStudy (anonymous):

Then the problem is written poorly. You probably have 4/27 on your answer guide, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The problem is definitely written poorly and you should tell your teacher, because they left out red FIRST and then blue. That is what they intended, but they didn't say or imply that.

OpenStudy (anonymous):

nope but thanks for helping me

OpenStudy (anonymous):

The way the problem reads, it could be the case that blue gets picked first and then red and then you would still have red and blue. Bad problem.

OpenStudy (anonymous):

yea i know

OpenStudy (anonymous):

I have a brother who has a doctorate in education, and he knows nothing about math. And it's people like him who design a lot of these questions!

OpenStudy (anonymous):

ha Probably :) can you help on my next one

OpenStudy (anonymous):

Sure, why not?